OpenStudy (anonymous):

Graph: 4x + 3y = -12

OpenStudy (compassionate):

Do you want me too solve for x or y?

OpenStudy (anonymous):

It just says to graph it, but I never done this type of math before, so I am confused.

OpenStudy (anonymous):

Start by isolating y So 4x+3y=-12 3y=-4x-12 y=-4/3x-12/3 y=-4/3x-4

OpenStudy (anonymous):

Any online grapher or graphing calculator should be able to do the rest for you.

OpenStudy (anonymous):

Basically -4 is where the line will cross through x=0 the slope would be down 4 to the right 3

OpenStudy (anonymous):

or a another way is to find the x and y intercepts (by setting x=0 solve for y, then set y=0, solve for x) then just draw a line connecting those two points...

OpenStudy (anonymous):

So I would use the y=mx+b method to solve it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

God, I hate math gives me such a headache.

OpenStudy (anonymous):

Just be patient with yourself and practice a lot. It does get easier.

OpenStudy (anonymous):

Thanks, I think it is because my teachers never took the time to explain and just kept passing me even with failing grades.

OpenStudy (anonymous):

Good luck!

OpenStudy (anonymous):

Thanks.